# Historically Relevant Event Structuring for Temporal Knowledge Graph Reasoning

### v1

@@ 4.4. Sensitivity Analysis (RQ5)
To further explore the sensitivity of $\mathsf{HisRES}$ in different settings (Q5), we conduct experiments to explore the influence of (a) granularity level and (b) hidden layers of GNNs on the performance of $\mathsf{HisRES}$, which is shown in Figure 5.
Figure 5. Sensitivity analysis on ICEWS14s dataset.
4.4.1. Exploring the Influence of Granularity Level
@@
The granularity level plays a pivotal role in controlling the interaction range inter-snapshot in the proposed multi-granularity evolutionary encoder simultaneously. It determines the quality of entity representations studied from the most recent snapshots. As illustrated in Figure 5(a), $\mathsf{HisRES}$ exhibits a compelling performance with capturing correlations at each 2 adjacent snapshots level. These findings highlight the impressive ability of $\mathsf{HisRES}$ to adapt to historical information from varying time spans. It is worth noting that $\mathsf{HisRES}$ maintains strong performance with different levels of granularity, demonstrating the robustness and effectiveness of our basic architecture.
4.4.2. Exploring the Influence of Hidden Layer of GNNs
@@
Since $\mathsf{HisRES}$ integrates CompGCN and ConvGAT to learn multi-granularity evolutionary and globally relevant patterns, respectively, it is essential to evaluate the hidden layer numbers by investigating the influence of aggregation multi-hop neighbor information on entity representation. The results are depicted in Figure 5(b). Specifically, the two-hop results outperform both the one-hop and three-hop results. This observation suggests that overly simplistic neighborhoods or increasing the number of hops does not significantly enhance the performance of aggregating relevant events. Conversely, it may lead to oversmoothing or introduce unnecessary features that impede effective representations.
5. Conclusion
@@
This paper introduces $\mathsf{HisRES}$, an innovative model for TKG reasoning. It comprises two novel encoders that excel at learning local and global entity representations by structuring historically relevant events. Specifically, we propose a multi-granularity evolutionary encoder to capture spatial-temporal interactions from inter-snapshot and intra-snapshot perspectives. In particular, we introduce a global relevance encoder featuring a distinctive ConvGAT to attend to crucial dependencies throughout the entire history. Furthermore, $\mathsf{HisRES}$ incorporates a self-gating mechanism for comprehensive fusion of TKG information. Experimental results on four event-based TKG datasets and various ablation studies demonstrate the effectiveness and superior performance of $\mathsf{HisRES}$. Future research avenues include exploring pruning techniques for global relevance and enhancing the modeling of temporal information.
Acknowledgements.

### v2

@@ IV-D Sensitivity Analysis (RQ5)
To further investigate the sensitivity of HisRES in different settings (Q5), we conduct experiments to explore the influence of granularity span (§IV-D1), hidden layers of GNNs (§IV-D2), and historical length (§IV-D3) on the performance of HisRES, which is shown in Figure 5, 6, and 7, respectively.
Figure 5: The influence of granularity spans from each single snapshot to every three adjacent snapshots.
IV-D1 Exploring the Influence of Granularity Span
@@
The granularity span $\omega$ plays a pivotal role in controlling the interaction range inter-snapshot in the proposed multi-granularity evolutionary encoder simultaneously. It determines the quality of entity representations studied from the most recent snapshots. As illustrated in Figure 5, HisRES exhibits a compelling performance with capturing correlations at each 2 adjacent snapshots level. These findings highlight the impressive ability of HisRES to adapt to historical information from varying time spans. It is worth noting that HisRES maintains strong performance with different levels of granularity, demonstrating the robustness and effectiveness of our basic architecture.
Figure 6: The influence of hidden layers, varying from one layer to four layers of ConvGAT.
IV-D2 Exploring the Influence of Hidden Layer of GNNs
@@
Since HisRES integrates CompGCN and ConvGAT to learn multi-granularity evolutionary and globally relevant patterns, respectively, it is essential to evaluate the hidden layer numbers by investigating the influence of aggregation multi-hop neighbor information on entity representation. The results are depicted in Figure 6. Specifically, the two-hop results outperform both the one-hop and three-hop results. This observation suggests that overly simplistic neighborhoods or increasing the number of hops does not significantly enhance the performance of aggregating relevant events. Conversely, it may lead to oversmoothing or introduce unnecessary features that impede effective representations.
Figure 7: The influence of historical length.
IV-D3 Exploring the Influence of Historical Length
The historical length $l$ determines the dependency degree of HisRES on recent history. Specifically, this value directly affects the performance of the multi-granularity evolutionary encoder and is expected to be smaller, as we already employ global relevance encoder to learn a wider range of historical knowledge. As illustrated in Figure 7, when the length of the adjacent history increases, the performance of HisRES exhibits less fluctuation. This observation further demonstrates the effectiveness of HisRES in considering historical information from both local and global perspectives.
TABLE VI: Detailed execution times of HisRES on ICEWS14s dataset. All times are reported in seconds per epoch.
Main Modules Components Details Training Time Inference Time
Multi-granularity Evolutionary Encoder Modeling intra-snapshot Structuring 7.99 0.74
Aggregation 24.22 1.98
Modeling inter-snapshot Structuring 8.07 0.76
Aggregation 22.77 1.85
Self-gating1 - 0.04 $<$0.01
Global Relevance Encoder Global Relevance Structuring - 96.04 10.42
ConvGAT - 28.76 4.49
Self-gating2 - - 0.06 $<$0.01
IV-E Evaluation on the Impact of Self-gating
To investigate the impact of self-gating in HisRES, we conduct weight analysis to examine how it adaptively merges: 1) multi-granular representations, and 2) representations from the multi-granularity evolutionary encoder and global relevance encoder. Notably, the values for each entity represent the mean across dimension $d$, and the values in 2) are further averaged from both raw and inverse queries.
Figure 8: Distribution of self-gating weights between intra- and inter-snapshot encoding on ICEWS14s. Lighter (darker) blue indicates higher (lower) intra-snapshot encoding weights.
IV-E1 Weight between intra- and inter-snapshot
As illustrated in Figure 8, in the first self-gating, most snapshots heavily rely on features modeled by inter-snapshot encoding during testing. The average weight assigned to intra-snapshot modeling is only 0.11. This observation demonstrates the crucial role of historical information from inter-snapshot modeling in TKG reasoning.
Figure 9: Distribution of self-gating weights between multi-granular and global encoding on ICEWS14s. Lighter (darker) blue indicates higher (lower) multi-granular encoding weights.
IV-E2 Weight between Multi-granular and Global History
As shown in Figure 9, HisRES adaptively balances multi-granularity and global representations through the second self-gating. The average weight assigned to multi-granularity modeling is 0.26, indicating that globally relevant historical information provides numerous effective cues. Consequently, both 1) and 2) demonstrate the effectiveness of self-gating in HisRES. Despite varying degrees of dependence on historical information across entities at different granularities, our model effectively learns appropriate representations from such diverse temporal granularities.
IV-F Execution Time Analysis
To evaluate the computational efficiency and performance trade-offs of HisRES, comprehensive timing analyses are conducted comparing our approach with SOTAs across diverse datasets, as illustrated in Figure 10. The datasets include ICEWS18 (higher entity count), ICEWS05-15 (extended time span), and GDELT (fine-grained timestamp and fact density), demonstrating the scalability of HisRES. Training and inference times are measured in seconds for consistent comparison. The observed variations in execution time can be attributed to three primary factors:
  1. 1.
    Variable optimal historical lengths significantly impact local encoding efficiency
  2. 2.
    Global information processing requires comprehensive fact consideration, especially the density of history, which may proportionally affect the efficiency
  3. 3.
    RETIA’s line graph construction introduces substantial computational overhead
The results demonstrate that HisRES maintains comparable computational efficiency. The predominant source of computational overhead in HisRES stems from global relevance structuring across the entire historical timeline, which is particularly sensitive to historical correlation density. For instance, ICEWS18 and GDELT exhibit high computational demands due to their dense fact distributions in global graphs. Conversely, ICEWS05-15, despite its large fact count, demonstrates lower density due to its extended temporal span. Notably, direct performance comparisons warrant careful interpretation due to methodological differences in data preprocessing. While existing approaches typically pre-compute and store historical graphs prior to training, HisRES dynamically generates multi-granular and globally relevant graphs during both training and inference phases. Although this real-time computation introduces additional overhead, it significantly enhances the model’s flexibility and practical applicability.
Figure 10: Execution time (seconds per epoch) comparison to baselines on four datasets.
All experiments are conducted using NVIDIA A800 80G GPUs, Intel Xeon Silver 4316 CPUs, and 256G RAM. Detailed analyses for the running time of each proposed module of HisRES are presented in Table VI. In the multi-granularity evolutionary encoder, the primary computational overhead stems from two processes: (1) structuring subgraphs at each granularity level (single snapshot or adjacent snapshots) and (2) aggregating information for each subgraph. Our experimental results demonstrate that structuring adjacent snapshots introduces only minimal additional processing time compared to structuring single snapshot subgraphs (8.07 seconds versus 7.99 seconds), with trivial differences in aggregation time (24.22 seconds versus 22.77 seconds). Although our computational analysis indicates that global relevance structuring constitutes a significant portion of HisRES’s total overhead, the efficiency of the proposed ConvGAT in processing numerous global historical facts is comparable to the aggregation time required for a single snapshot. Additionally, our results confirm that the self-gating mechanism represents a highly cost-effective module, efficiently balancing weights across different modules with an inference time of less than 0.01 seconds.
TABLE VII: Case study of HisRES using three queries from the ICEWS14s dataset, dated 2014-12-01. The ground truth entity is shown in bold. † denotes attention scores derived from non-query relations. - indicates the objects are not in $\mathcal{G}^{\mathcal{H}}_{2014-12-01}$.
Query Hits@5 of HisRES Attention Score from ConvGAT Truth in History
s: Philippines (a) r: Arrest,_detain,_or_charge_with_legal_action t: 2014-12-01 Number of links in $\mathcal{G}^{\mathcal{H}}_{2014-12-01}$: 47 Men_(Philippines) 0.0030 2014-01-17
Criminal_(Philippines) (d) 0.0307 2014-02-06
Abu_Sayyaf 0.0026 2014-05-30
Employee_(Philippines) 0.0025 2014-08-27
Businessperson_(Philippines) 0.0020 ……
s: Police_(Philippines) (b) r: Confiscate_property t: 2014-12-01 Number of links in $\mathcal{G}^{\mathcal{H}}_{2014-12-01}$: 8 Criminal_(Philippines) (d) 0.1126 2014-02-21
Philippines 0.3459† 2014-05-22
Abu_Sayyaf 0.0790 2014-06-24
Men_(Philippines) - 2014-07-13
Employee_(Philippines) - ……
s: City_Mayor_(Philippines) (c) r: Make_statement t: 2014-12-01 Number of links in $\mathcal{G}^{\mathcal{H}}_{2014-12-01}$: 16 Philippines 0.3632†
City_Mayor_(Philippines) - 2014-04-09
Police_(Philippines) 0.0933 2014-09-11
Military_(Philippines) 0.0261 2014-09-18
Criminal_(Philippines) (d) 0.0957
IV-G Robustness Analysis
To investigate the robustness of HisRES, we evaluated performance by introducing random Gaussian noise of varying intensities to the initial entity embeddings during testing. This evaluation is conducted on HisRES, TiRGN [7], and RE-GCN [6], with results denoted by the ”-noise” suffix in Figure 11. Experimental results demonstrate that HisRES maintains robust performance under various noise intensities. This resilience stems from HisRES’s global relevance encoder, which effectively identifies highly correlated facts in the global history. Quantitatively, on ICEWS14s, HisRES shows moderate noise resistance with a 17.79% performance degradation at noise intensity 0.15, compared to degradations of 20.60% and 49.49% for TiRGN and RE-GCN, respectively. Though TiRGN’s historical information mask matrix helps mitigate performance decline under noise, HisRES maintains better performance for noise intensities below 0.11 on ICEWS18.
Figure 11: Analysis of MRR performance under varying Gaussian noise intensities on ICEWS14s and ICEWS18. Light blue, green, and purple shading indicate the variation range of HisRES, TiRGN, and RE-GCN, respectively.
IV-H Case Study of the proposed ConvGAT
To demonstrate the capability of HisRES to focus on crucial events within a global historical context and to showcase the effectiveness of our proposed ConvGAT, we select three queries pertaining to Philippines from ICEWS14s, all dated 2014-12-01, as illustrative examples:
  1. 1.
    (Philippines, Arrest,_detain,_or_charge_with_legal_action, ? , 2014-12-01);
  2. 2.
    (Police_(Philippines), Confiscate_property, ? , 2014-12-01);
  3. 3.
    (City_Mayor_(Philippines), Make_statement, ? , 2014-12-01).
All three queries share the ground truth entity Criminal_(Philippines). Table VII presents comprehensive information for each query, including: the number of incoming edges for each query’s subject in the globally relevant graph $\mathcal{G}^{\mathcal{H}}_{2014-12-01}$, the Hits@5 objects predicted by HisRES, the attention scores calculated by ConvGAT for predicted objects in $\mathcal{G}^{\mathcal{H}}_{2014-12-01}$, and the historical timestamps when the ground truth facts occurred. For readability, we denote the subjects and ground truth as a to d. These examples, derived from historical events, exhibit varying densities of association facts in $\mathcal{G}^{\mathcal{H}}_{2014-12-01}$ (excluding inverse queries). ConvGAT demonstrates its ability to identify crucial facts from $\mathcal{G}^{\mathcal{H}}_{2014-12-01}$, regardless of their temporal distance, and to effectively aggregate information from entities highly related to Philippines.
Table VII reveals that, for all three queries, ConvGAT assigns the highest attention scores to the ground truth d under the respective query relations. Specifically, in example 1), despite 47 entities being directly relevant to a, HisRES successfully identifies and prioritizes important entities with high predicted scores, even when they are not from the most recent events. As shown in the first part of Table VII, d contributes the highest score of 0.0307. For example 2), d provides the highest attention to b, with a score of 0.1126. Furthermore, ConvGAT effectively aggregates related entities (e.g., a) under other relations with a cumulative attention score of 0.3459, demonstrating robust semantic relevance. In example 3), although the ground truth is ranked at a lower position, ConvGAT nonetheless ensures that its global attention score relative to query remains high.
V Conclusion
@@
This paper introduces HisRES, an innovative model for TKG reasoning. It comprises two novel encoders that excel at learning local and global representations by structuring historically relevant events. Specifically, we propose a multi-granularity evolutionary encoder to capture structural and temporal interactions from intra-snapshot and inter-snapshot perspectives. In particular, we introduce a global relevance encoder featuring a distinctive ConvGAT to attend to crucial dependencies throughout the entire history. Furthermore, HisRES incorporates a self-gating mechanism for comprehensive fusion of TKG information. Experimental results on four event-based TKG datasets and various ablation studies demonstrate the effectiveness and superior performance of HisRES. Future research avenues include exploring pruning techniques for global relevance and enhancing the modeling of temporal information.
Acknowledgments
This research work is partly supported by National Natural Science Foundation of China [No. 62376055].